Find the derivative of the function



P(x)=ln ⁡[ (4x + 1)^3 / (2x − 5)^4 ]


is


a. −4(2x−17) / (4x+1)(2x−5)

b.−4(2x−17) / (4x+1)

c. 4(−2x−17) / (4x+1)(2x−5)

d. (−2x−17) / (4x+1)(2x−5)



1
Expert's answer
2022-04-25T18:00:46-0400

"P\u2019(x)=(ln \u2061[ (4x + 1)^3 \/ (2x \u2212 5)^4 ])\u2019= \\frac{1}{(4x + 1)^3 \/ (2x \u2212 5)^4}*((4x + 1)^3 \/ (2x \u2212 5)^4)\u2019=\\frac{(2x-5)^4}{(4x+1)^3}*\\frac{12(4x+1)^2*(2x-5)^4-8(2x-5)^3*(4x+1)^3}{(2x-5)^8}=\\frac{12(2x-5)-8(4x+1)}{(4x+1)(2x-5)}=\\frac{24x-60-32x-8}{(4x+1)(2x-5)}=\\frac{4(-2x-17)}{(4x+1)(2x-5)}"